# Conversational LLMs Simplify Secure Clinical Data Access, Understanding, and Analysis

### v1

@@ Abstract
As ever-larger clinical datasets become available, they have the potential to unlock unprecedented opportunities for medical research. Foremost among them is Medical Information Mart for Intensive Care (MIMIC-IV), the world’s largest open-source EHR database. However, the inherent complexity of these datasets, particularly the need for sophisticated querying skills and the need to understand the underlying clinical settings, often presents a significant barrier to their effective use. M3 lowers the technical barrier to understanding and querying MIMIC-IV data. With a single command it retrieves MIMIC-IV from PhysioNet, launches a local SQLite instance (or hooks into the hosted BigQuery), and-via the Model Context Protocol (MCP)-lets researchers converse with the database in plain English. Ask a clinical question in natural language; M3 uses a language model to translate it into SQL, executes the query against the MIMIC-IV dataset, and returns structured results alongside the underlying query for verifiability and reproducibility. Demonstrations show that minutes of dialogue with M3 yield the kind of nuanced cohort analyses that once demanded hours of handcrafted SQL and relied on understanding the complexities of clinical workflows. By simplifying access, M3 invites the broader research community to mine clinical critical-care data and accelerates the translation of raw records into actionable insight.
@@ 1.1 The Challenge of Analyzing Large-Scale Clinical Databases
The digital transformation of healthcare has led to the generation and accumulation of vast quantities of electronic health record (EHR) data [1], creating invaluable resources for secondary use of data, such as medical research that offer deep insights into disease patterns, treatment efficacy, and patient outcomes. However, the barrier to use these datasets is often high, due to data inherent complexity and required data querying skills. More in detail, clinical databases are typically relational, consisting of numerous interconnected tables and a multitude of fields, each with specific definitions and coding schemes. Effectively navigating and extracting meaningful information from such intricate structures necessitates specialized technical skills, primarily proficiency in SQL, a thorough understanding of the database schema, and how data points are temporarily connected and semantically linked. The SQL technical requirement forms a substantial barrier for many clinical researchers, while the clinical data domain poses a barrier to entry to most non-clinical data scientists or analysts. Consequently, the technical skill set needed to directly query complex databases like MIMIC-IV [2] can limit the pool of researchers able to leverage these resources, potentially impeding the pace of innovation (e.g., clinical process improvement). This also highlights an interdisciplinary gap where clinical experts, who formulate the critical research questions, may be disconnected from the data extraction process, which often falls to data scientists or programmers. Tools that can bridge this divide by simplifying data access are therefore of growing importance, and some are already in use in academic medical centers and other clinical settings. [3, 4, 5]
Anthropic’s Model Context Protocol (MCP) [6] provides a standardized framework for managing AI model interactions with external software tools and data sources, offering a promising approach to address these accessibility challenges through secure and controlled interfaces.
@@ 1.2 The Role of MIMIC-IV in Critical Care Research
MIMIC-IV stands as a cornerstone of publicly available database for critical care research. Developed by the MIT Laboratory for Computational Physiology, this dataset contains de-identified health data associated with patients admitted to intensive care units (ICUs) or the emergency department at the Beth Israel Deaconess Medical Center. MIMIC-IV (version 3.1) [7] includes data from approximately 364,627 unique individuals (each represented by a unique subject_id), 546,028 hospitalizations and 94,458 unique ICU stays. The dataset is rich in detail, including patient demographics, vital sign measurements, laboratory test results, medications, procedures, and more.
MIMIC-IV is widely utilized in the research community for developing and validating clinical prediction models, understanding disease trajectories, evaluating treatment interventions, and ultimately aiming to improve patient care in critical settings. The availability of MIMIC-IV through the PhysioNet platform [7], which provides access modalities such as Google BigQuery for the full dataset, enhances research transparency and reproducibility, key elements of scientific progress [8]. The public, albeit credentialed, nature of MIMIC-IV enabled numerous research groups to work with standardized, high-fidelity clinical data, fostering collaboration and building upon prior work.
@@ 1.3 Introducing M3: Objectives and Contributions
We evaluate M3 using samples from the EHRSQL 2024 test set [9], a benchmark designed for assessing natural language-to-SQL performance in clinical contexts. This dataset is based on the publicly available MIMIC-IV demo (v2.2) [10], which includes a representative subset of real hospital data and is freely accessible through PhysioNet. To illustrate the type of complex question that M3 can handle, Figure 1 shows the result of a query involving multiple temporal and clinical constraints. Specifically, this complex request inquires about the top three most commonly prescribed medications that were administered after a diagnosis of unspecified anemia (ICD code) during the same hospital admission, among patients aged 60 or older whose diagnosis occurred in the year 2100 or later. This result was obtained through the M3 system powered by Claude Sonnet 4 via the MCP. The example highlights how M3 enables non-technical users to retrieve clinically meaningful insights from complex databases using only natural language.
@@
Listing 1: Correct SQL query [9]
@@ 2.1 Evolution of Clinical Database Access Tools
Recent years have seen significant progress in lowering the technical barriers to accessing and analyzing complex clinical databases, particularly for researchers without advanced programming expertise. Early efforts focused on direct SQL querying and basic graphical interfaces, requiring significant technical expertise from users. The MIMIC-II project [11] introduced web-based query builders and virtual machine environments, marking an important step toward simplifying database access for clinical researchers.
The development of MIMIC-IV expanded these capabilities through various access modalities, including cloud platforms such as Google BigQuery [12, 2]. While this improved data accessibility and processing capabilities, the fundamental challenge of SQL expertise remained and was always compounded by the equally important required understanding of the clinical domain. Visual query builders and curated SQL templates [13] have attempted to bridge this gap, though often sacrificing query flexibility for ease of use.
The emergence of standards such as HL7 FHIR, the OMOP Common Data Model, and mCODE is enabling new, more scalable methods of accessing and sharing health data. The MIMIC-IV on FHIR implementation represents an important step toward standardized data access, though it brings its own complexities in terms of resource modeling and query patterns [14, 15, 16] .
@@ 2.2 Natural Language Interfaces for Medical Data
The development of natural language interfaces for databases (NLIDB) has seen several approaches evolve in parallel. Early NLIDB implementations on healthcare domain like MIMICSQL [17] demonstrated the basic feasibility of translating natural language to SQL, though they often struggled with query complexity and medical terminology variations. Subsequent systems such as EHRSQL [18] employed more sophisticated techniques to improve query understanding, showing better handling of medical terminology while still facing challenges with complex temporal relationships and nested queries common in clinical research.
@@ 2.3 Benchmarks and Evaluation Frameworks
The development of specialized benchmarks has been crucial for advancing the field. While general text-to-SQL benchmarks like BIRD [19], Spider [20] and WikiSQL [21] provided foundational evaluation frameworks, they lack medical domain coverage and specificity. More recent efforts such as BiomedSQL [22] and the EHRSQL 2024 shared task [23] have introduced domain-specific challenges that better reflect real-world clinical querying needs. These benchmarks have revealed significant challenges in handling implicit medical knowledge, understanding temporal relationships in clinical data, managing hierarchical medical concepts, and integration with clinical workflows.
@@ 2.4 Security and Integration Frameworks
Security considerations in clinical database access have evolved from basic database-level security and input sanitization, as outlined in resources like the OWASP SQL Injection Prevention Cheat Sheet [24], to more comprehensive approaches. The introduction of the MCP  [6] represents a significant advance in AI-database integration that can support modern security standards, providing precise interaction patterns, access control mechanisms, audit capabilities, and reproducible query execution. Industry adoption of MCP has indeed grown across various domains including software development, scientific research, and biomedical [25], [26].
@@ 2.5 Current Challenges and Opportunities
Existing solutions continue to face several key challenges. General-purpose text-to-SQL systems often struggle with medical terminology and relationships, while specialized medical systems may sacrifice query flexibility for security. Many current solutions lack robust mechanisms for ensuring query provenance and result reproduction. Technical integration requirements can remain substantial, and scaling to handle the complexity of full clinical databases presents ongoing challenges. To our knowledge, none of these is currently integrated in a desktop generative AI application, such as Claude.
@@ 3.2 Data Sources and Access Layer
  • Local SQLite Database: For rapid prototyping and development, M3 provides a local SQLite implementation using the official 100-patient demo subset of MIMIC-IV [2]. This option requires minimal setup and incurs no cloud costs. The system handles the complete Extract, Transform, Load (ETL) process from PhysioNet data files to a local database, including schema inference and standardized null value handling.
  • Google BigQuery: For full-scale research, M3 connects to the complete MIMIC-IV v3.1 schemas [7] on Google BigQuery. This implementation supports advanced features such as parameterized queries, cost estimation, and IAM-based access control. Access requires prior PhysioNet credentialing and an active Google Cloud project.
@@ 3.4.1 Core Database Tools
The foundational layer provides essential database access capabilities including schema discovery, table inspection, and query execution. These tools allow agents to understand the database structure and execute validated SQL queries against the MIMIC-IV dataset.
@@ 4 Results
To assess the capabilities of the M3 system, we performed an evaluation using the challenging EHRSQL 2024 benchmark [23]. This benchmark is a prominent and specialized challenge for assessing the performance of text-to-SQL systems on clinical data, using the MIMIC-IV demo database [10]. Our goal was to measure the system’s accuracy in a realistic setting and to understand the qualitative nature of its successes and failures.
@@ 4.1 Evaluation Methodology
The experimental setup consisted of the M3 system powered by Claude Sonnet 4 through the Model Context Protocol. We utilized the ‘mimic_iv.sqlite‘ database, which is the official database for the EHRSQL task and is based on the MIMIC-IV demo version 2.2 [12]. This ensures that our results are directly comparable to the context of the EHRSQL benchmark.
The official EHRSQL benchmark code defines a fixed ’current time’ of "2100-12-31 23:59:00" for evaluating temporal queries [27]. To align our independent M3 system with this requirement, we simulated the condition by adding a contextual instruction to the start of each relevant prompt: "Set the current time to be “2100-12-31 23:59:00” when using m3 mcp." This step was essential for faithfully replicating the benchmark’s environment and validating our results.
@@ 4.2 Quantitative Performance
Out of the 100 answerable questions, the M3 system correctly generated the SQL and provided the right answer for 94 of them, yielding a simple execution accuracy of 94%. A detailed breakdown of the performance is presented in Table 1.
@@
Outcome Count
Correct Answers 94
Incorrect Answers 6
   – Due to Linguistic Ambiguity 1
   – Due to Logical/Execution Errors 5
Total Evaluated 100
The reported 94% accuracy was determined through a meticulous human evaluation process. For each of the 100 questions, the final answer generated by the M3 system was manually reviewed and compared against the ground truth answer from the EHRSQL dataset. An answer was deemed correct if it was logically and semantically equivalent to the ground truth, even if the phrasing or presentation differed. This reliance on human judgment is a necessary and standard practice for evaluating complex question-answering systems, as automated scripts can fail to capture the correctness of varied but logically sound responses. This evaluation approach is consistent with methodologies used in the development of other large-scale text-to-SQL benchmarks [28].
@@ 4.3 Visual Examples of Complex Query Results
To complement the quantitative evaluation, we also present several illustrative examples of complex queries processed by M3 on MIMIC-IV demo [10], together with their corresponding visualized outputs (Figures  3 and  4). These were generated using the MIMIC-IV demo database via the Claude-powered M3 system. Each example includes the natural language query and the resulting visualisation, designed to reflect real-world clinical insights extractable from MIMIC-IV.
@@
Figure 4: Query: “Among sepsis patients, what’s the source-of-infection distribution and how do groups differ in ICU stay and mortality?”
@@
4.4 Qualitative Error Analysis
A manual inspection of the six incorrect responses provided critical insights into the system’s limitations. We found that only one of the errors was caused by ambiguous phrasing in the natural language question, while the remaining five were due to various types of SQL generation mistakes.
A clear example of the ambiguity issue was a query asking: "How many days have passed since patient X’s last stay in the discharge lounge?" The model interpreted this as calculating time since the end of the stay (exit time), while the ground truth query calculated from the start (entry time). Since both interpretations are clinically plausible and the question did not specify a reference point, this discrepancy highlights a true instance of linguistic ambiguity. This type of ambiguity is a significant challenge in clinical NLP: the same phrase can have multiple valid interpretations [29], and the correct one often depends on unstated clinical conventions that generalized models may lack [30].
The remaining five errors were attributed to diverse failure modes in SQL generation, including incorrect filtering, misinterpretation of temporal or logical constraints, and oversights in query construction. For example, one error involved counting patients as "current" based on survival rather than ongoing hospitalization, and another omitted prescribed drugs that met the temporal criteria but were unrelated to detoxification. These errors represent more traditional text-to-SQL challenges, where the complexity of the database schema or the question leads to incorrect query structure [20].
While only one error was formally classified as due to ambiguity, several other failures might plausibly have been avoided if the questions had been phrased with greater precision. For instance, more explicit wording around "difference in blood pressure measurements" or "all drugs prescribed after detoxification" might have helped the model interpret the intent more accurately. This observation aligns with a broader trend in clinical NLP: even when a model is technically capable, subtle underspecification or lack of domain-specific language norms in a question can lead to incorrect outputs. Improving question clarity—especially in benchmarking contexts—could improve system performance and reduce apparent errors that are not fundamentally due to model capability but to task formulation.

### v3

# Renato Umeton1,4, Amelia Fiske 2, Leo Anthony Celi1,5

@@ Abstract
Large-scale clinical databases offer opportunities for medical research, but their complexity creates barriers to effective use. The Medical Information Mart for Intensive Care (MIMIC-IV), one of the world’s largest open-source electronic health record databases, traditionally requires both SQL proficiency and clinical domain expertise. We introduce M3, a system that enables natural language querying of MIMIC-IV data through the Model Context Protocol. With a single command, M3 retrieves MIMIC-IV from PhysioNet, launches a local SQLite instance or connects to hosted BigQuery, and allows researchers to pose clinical questions in plain English. We evaluated M3 using one hundred questions from the EHRSQL 2024 benchmark with two language models: the proprietary Claude Sonnet 4 achieved 94% accuracy, while the open-source gpt-oss-20B (deployable locally on consumer hardware) achieved 93% accuracy. Both models translate natural language into SQL, execute queries against MIMIC-IV, and return structured results alongside the underlying query for verification. Error analysis revealed that most failures stemmed from complex temporal reasoning or ambiguous question phrasing rather than fundamental architectural limitations. The comparable performance of a smaller open-source model demonstrates that privacy-preserving local deployment is viable for sensitive clinical data analysis. M3 lowers technical barriers to critical care data analysis while maintaining security through OAuth2 authentication, query validation, and comprehensive audit logging.
@@ 1.1 The Challenge of Analyzing Large-Scale Clinical Databases
The digital transformation of healthcare has led to the generation and accumulation of vast quantities of electronic health record (EHR) data [evans2016electronic], creating invaluable resources for secondary use of data, such as medical research that offer deep insights into disease patterns, treatment efficacy, and patient outcomes. However, the barrier to use these datasets is often high, due to data inherent complexity and required data querying skills. More in detail, clinical databases are typically relational, consisting of numerous interconnected tables and a multitude of fields, each with specific definitions and coding schemes. Effectively navigating and extracting meaningful information from such intricate structures necessitates specialized technical skills, primarily proficiency in SQL, a thorough understanding of the database schema, and how data points are temporarily connected and semantically linked. The SQL technical requirement forms a substantial barrier for many clinical researchers, while the clinical data domain poses a barrier to entry to most non-clinical data scientists or analysts. Consequently, the technical skill set needed to directly query complex databases like MIMIC-IV [johnson2023mimiciv] can limit the pool of researchers able to leverage these resources, potentially impeding the pace of innovation (e.g., clinical process improvement). This also highlights an interdisciplinary gap where clinical experts, who formulate the critical research questions, may be disconnected from the data extraction process, which often falls to data scientists or programmers. Tools that can bridge this divide by simplifying data access are therefore of growing importance, and some are already in use in academic medical centers and other clinical settings. [statnewsSvelte, jiang2023health, doi:10.1056/AIcs2300191]
Anthropic’s Model Context Protocol (MCP) [anthropic2024mcp] provides a standardized framework for managing AI model interactions with external software tools and data sources, offering a promising approach to address these accessibility challenges through secure and controlled interfaces.
@@ 1.2 The Role of MIMIC-IV in Critical Care Research
MIMIC-IV stands as a cornerstone of publicly available database for critical care research. Developed by the MIT Laboratory for Computational Physiology, this dataset contains de-identified health data associated with patients admitted to intensive care units (ICUs) or the emergency department at the Beth Israel Deaconess Medical Center. MIMIC-IV (version 3.1) [mimicivphysionet] includes data from approximately 364,627 unique individuals (each represented by a unique subject_id), 546,028 hospitalizations and 94,458 unique ICU stays. The dataset is rich in detail, including patient demographics, vital sign measurements, laboratory test results, medications, procedures, and more.
MIMIC-IV is widely utilized in the research community for developing and validating clinical prediction models, understanding disease trajectories, evaluating treatment interventions, and ultimately aiming to improve patient care in critical settings. The availability of MIMIC-IV through the PhysioNet platform [mimicivphysionet], which provides access modalities such as Google BigQuery for the full dataset, enhances research transparency and reproducibility, key elements of scientific progress [goldberger2000physionet]. The public, albeit credentialed, nature of MIMIC-IV enabled numerous research groups to work with standardized, high-fidelity clinical data, fostering collaboration and building upon prior work.
@@ 1.3 Introducing M3: Objectives and Contributions
We evaluate M3 using samples from the EHRSQL 2024 test set [ehrsql2024github], a benchmark designed for assessing natural language-to-SQL performance in clinical contexts. This dataset is based on the publicly available MIMIC-IV demo (v2.2) [johnson2023mimiciv_demo], which includes a representative subset of real hospital data and is freely accessible through PhysioNet. To illustrate the type of complex question that M3 can handle, Figure 1 shows the result of a query involving multiple temporal and clinical constraints. Specifically, this complex request inquires about the top three most commonly prescribed medications that were administered after a diagnosis of unspecified anemia (ICD code) during the same hospital admission, among patients aged 60 or older whose diagnosis occurred in the year 2100 or later. This result was obtained through the M3 system powered by Claude Sonnet 4  [anthropic2025claude4] via the MCP. The example highlights how M3 enables non-technical users to retrieve clinically meaningful insights from complex databases using only natural language.
@@
Listing 1: Correct SQL query [ehrsql2024github]
@@
We also include a dedicated ethical considerations section  5 to reflect on the broader implications of lowering access barriers to clinical data via AI systems.
@@ 2.1 Evolution of Clinical Database Access Tools
Recent years have seen significant progress in lowering the technical barriers to accessing and analyzing complex clinical databases, particularly for researchers without advanced programming expertise. Early efforts focused on direct SQL querying and basic graphical interfaces, requiring significant technical expertise from users. The MIMIC-II project [saeed2011mimicii] introduced web-based query builders and virtual machine environments, marking an important step toward simplifying database access for clinical researchers.
The development of MIMIC-IV expanded these capabilities through various access modalities, including cloud platforms such as Google BigQuery [johnson2022mimiciv, johnson2023mimiciv]. While this improved data accessibility and processing capabilities, the fundamental challenge of SQL expertise remained and was always compounded by the equally important required understanding of the clinical domain. Visual query builders and curated SQL templates [tian2024sqlucid] have attempted to bridge this gap, though often sacrificing query flexibility for ease of use.
The emergence of standards such as HL7 FHIR, the OMOP Common Data Model, and mCODE is enabling new, more scalable methods of accessing and sharing health data. The MIMIC-IV on FHIR implementation represents an important step toward standardized data access, though it brings its own complexities in terms of resource modeling and query patterns [hl7fhir, bennett2023mimic, ohdsiStandardization] .
@@ 2.2 Natural Language Interfaces for Medical Data
The development of natural language interfaces for databases (NLIDB) has seen several approaches evolve in parallel. Early NLIDB implementations on healthcare domain like MIMICSQL [wang2020text] demonstrated the basic feasibility of translating natural language to SQL, though they often struggled with query complexity and medical terminology variations. Subsequent systems such as EHRSQL [yu2023ehrsql] employed more sophisticated techniques to improve query understanding, showing better handling of medical terminology while still facing challenges with complex temporal relationships and nested queries common in clinical research.
@@ 2.3 Benchmarks and Evaluation Frameworks
The development of specialized benchmarks has been crucial for advancing the field. While general text-to-SQL benchmarks like BIRD [li2024can], Spider [yu2018spider] and WikiSQL [zhong2017seq2sql] provided foundational evaluation frameworks, they lack medical domain coverage and specificity. More recent efforts such as BiomedSQL [koretsky2025biomedsql] and the EHRSQL 2024 shared task [choi2024ehrsql] have introduced domain-specific challenges that better reflect real-world clinical querying needs. These benchmarks have revealed significant challenges in handling implicit medical knowledge, understanding temporal relationships in clinical data, managing hierarchical medical concepts, and integration with clinical workflows.
@@ 2.4 Security and Integration Frameworks
Security considerations in clinical database access have evolved from basic database-level security and input sanitization, as outlined in resources like the OWASP SQL Injection Prevention Cheat Sheet [owasp2023sqlcheat], to more comprehensive approaches. The introduction of the MCP  [anthropic2024mcp] represents a significant advance in AI-database integration that can support modern security standards, providing precise interaction patterns, access control mechanisms, audit capabilities, and reproducible query execution. Industry adoption of MCP has indeed grown across various domains including software development, scientific research, and biomedical [mcpharma2025], [mcpsuperagi2025].
@@ 2.5 Current Challenges and Opportunities
Existing solutions continue to face several key challenges. General-purpose text-to-SQL systems often struggle with medical terminology and relationships, while specialized medical systems may sacrifice query flexibility for security. Many current solutions lack robust mechanisms for ensuring query provenance and result reproduction. Technical integration requirements can remain substantial, and scaling to handle the complexity of full clinical databases presents ongoing challenges. To our knowledge, none of these is currently integrated in a desktop generative AI application, such as Claude Desktop [anthropic2025desktop] for instance.
@@ 3.2 Data Sources and Access Layer
  • Local SQLite Database: For rapid prototyping and development, M3 provides a local SQLite implementation using the official 100-patient demo subset of MIMIC-IV [johnson2023mimiciv]. This option requires minimal setup and incurs no cloud costs. The system handles the complete Extract, Transform, Load (ETL) process from PhysioNet data files to a local database, including schema inference and standardized null value handling.
  • Google BigQuery: For full-scale research, M3 connects to the complete MIMIC-IV v3.1 schemas [mimicivphysionet] on Google BigQuery. This implementation supports advanced features such as parameterized queries, cost estimation, and IAM-based access control. Access requires prior PhysioNet credentialing and an active Google Cloud project.
@@ 3.4.1 Core Database Tools
The foundational layer provides essential database access capabilities including schema discovery, table inspection, and query execution. These tools allow agents to understand the database structure and execute flexible and transparent SQL queries against the MIMIC-IV dataset.
@@ 4 Results
To assess the capabilities of the M3 system, we performed an evaluation using the challenging EHRSQL 2024 benchmark [choi2024ehrsql]. This benchmark is a prominent and specialized challenge for assessing the performance of text-to-SQL systems on clinical data, using the MIMIC-IV demo database [johnson2023mimiciv_demo]. Our goal was to measure the system’s accuracy in a realistic setting and to understand the qualitative nature of its successes and failures.
@@ 4.1 Evaluation Methodology
The experimental setup consisted of M3 powered by two different language models: Claude Sonnet 4 (a proprietary frontier model) [anthropic2025claude4] and gpt-oss-20B [openai_gpt_oss_20b] (an open-source model). We utilized the ‘mimic_iv.sqlite‘ database from the EHRSQL 2024 benchmark repository [ehrsql_github], which is based on MIMIC-IV demo version 2.2 [johnson2022mimiciv]. For the open-source model evaluation, we used LM Studio [lmstudio] to host gpt-oss-20B and ran tests on a MacBook M1 Max with 32GB RAM, demonstrating feasibility of local deployment on consumer hardware. We utilized the ‘mimic_iv.sqlite‘ database, which is the official database for the EHRSQL task and is based on the MIMIC-IV demo version 2.2 [johnson2022mimiciv]. This ensures that our results are directly comparable to the context of the EHRSQL benchmark.
The official EHRSQL benchmark code defines a fixed ’current time’ of "2100-12-31 23:59:00" for evaluating temporal queries [ehrsql_github]. To align our independent M3 system with this requirement, we simulated the condition by adding a contextual instruction to the start of each relevant prompt: "Set the current time to be “2100-12-31 23:59:00” when using m3 mcp." This step was essential for faithfully replicating the benchmark’s environment and validating our results.
@@ 4.2 Quantitative Performance
Out of one hundred answerable questions, both models demonstrated high accuracy. Claude Sonnet 4 correctly generated appropriate SQL queries and provided the right answer for 94 questions (94% accuracy), whereas gpt-oss-20B achieved 93% accuracy.
The comparable performance between Claude Sonnet 4 and the smaller open-source gpt-oss-20B is notable. gpt-oss-20B can be deployed locally on consumer hardware, offering a viable alternative for researchers with strict data privacy requirements or limited cloud connectivity. A detailed breakdown of the performance is presented in Table 1.
@@
Outcome Claude Sonnet 4 gpt-oss-20B
Correct Answers 94 93
Incorrect Answers 6 7
Total Evaluated 100 100
The reported 94% accuracy was determined through a meticulous human evaluation process. For each of the 100 questions, the final answer generated by the M3 system was manually reviewed and compared against the ground truth answer from the EHRSQL dataset. An answer was deemed correct if it was logically and semantically equivalent to the ground truth, even if the phrasing or presentation differed. For example, when the ground truth is “1" (yes), automated evaluation would incorrectly penalize a response of “Yes, the patient meets this criterion" despite being semantically equivalent. This reliance on human judgment is a necessary and standard practice for evaluating complex question-answering systems, as automated scripts can fail to capture the correctness of varied but logically sound responses. This evaluation approach is consistent with methodologies used in the development of other large-scale text-to-SQL benchmarks [pourreza2023evaluating].
@@ 4.3 Visual Examples of Complex Query Results
To complement the quantitative evaluation, we also present several illustrative examples of complex queries processed by M3 on MIMIC-IV demo [johnson2023mimiciv_demo], together with their corresponding visualized outputs (Figures  3 and  4). These were generated using the MIMIC-IV demo database via the Claude-powered M3 system. Each example includes the natural language query and the resulting visualisation, designed to reflect real-world clinical insights extractable from MIMIC-IV.
@@
Figure 4: Query: “Among sepsis patients, what’s the source‑of‑infection distribution and how do groups differ in ICU stay and mortality?”
@@
4.4 Model Comparison and Error Analysis
Analysis of the incorrect responses reveals both shared failure modes and model-specific weaknesses. Understanding these patterns provides insight into the limitations of current natural language-to-SQL systems in clinical contexts.
Common failures across both models. Three questions challenged both Claude Sonnet 4 and gpt-oss-20B, suggesting these questions contain inherent ambiguities or make unstated assumptions.
One question asked about days elapsed since a patient’s last discharge lounge stay. The gold query calculated from entry time (intime), while both models used exit time (outtime). The phrase “days since last stay” does not explicitly specify whether the reference point is the beginning or end of the stay, making both interpretations defensible.
Another question about the most frequent microbiology tests for hemodialysis patients did not specify the temporal relationship between tests and procedure. The gold query required tests to occur after hemodialysis during the same admission, but this ordering was not explicit in the question wording. Both models returned different test lists, with some overlap. The discrepancy appears to stem from how each model handled cases where multiple tests had identical frequencies, leading to different selections among equally-ranked options.
A third question asked about patients receiving serology/blood microbiology tests since 2100. Both models overcounted substantially (Claude: 41 patients, gpt-oss-20B: 40 patients) versus the gold standard (8 patients). The gold query searched for the exact specimen type ‘serology/blood’, while both models interpreted this phrase more broadly to include various blood-related tests. Without explicit enumeration of valid specimen type values, this broader interpretation represents a reasonable approach to an underspecified query.
Claude Sonnet 4-specific errors. Claude made three unique errors. A question asking for the difference between the last and second-to-last blood pressure measurements was incorrectly interpreted, with Claude comparing the first and second-to-last values instead, yielding +8.0 mmHg instead of the correct -9 mmHg. This appears to be a temporal ordering mistake in query construction.
A question about drugs prescribed after alcohol detoxification during the same hospital visit resulted in Claude identifying only phenobarbital-related medications while omitting other concurrent prescriptions (docusate sodium, haloperidol, latanoprost, omeprazole, sarna lotion). The model appears to have filtered for withdrawal-related drugs, though the question did not specify this restriction.
For a question asking for the count of current patients aged 60 or above, Claude returned 44 patients by counting those without a death date, while the correct interpretation required checking for ongoing hospitalization (dischtime IS NULL), which would yield 1 patient. This reflects the need for more precise terminology, distinguishing “currently admitted” from “alive” (not deceased) to avoid ambiguity in database queries.
gpt-oss-20B-specific errors. gpt-oss-20B made four unique errors. For a question about diagnoses following BMI 35.0-35.9 diagnosis, the model used ICD-9 code 278.00 (general obesity) as a proxy, explaining that BMI data was not directly available. However, the database actually contained a specific ICD code for “body mass index 35.0-35.9, adult” that the model failed to discover through schema exploration.
A question about patients receiving nutritional substance introduction after postprocedural pneumothorax resulted in gpt-oss-20B returning 0 patients while the ground truth indicated 1. This likely resulted from using hardcoded ICD codes rather than the gold query’s flexible approach of matching long title descriptions.
For a question asking for the first specimen test given to a patient since March 2100, gpt-oss-20B incorrectly identified “pt” (prothrombin time) from the labevents table, while the correct answer “mrsa screen” resided in the microbiologyevents table. This indicates confusion between laboratory test results and specimen collection procedures, which are tracked in different tables.
Finally, for a question asking for the top four most frequent lab tests, gpt-oss-20B’s results (glucose, chloride, sodium, hemoglobin) differed from the gold standard (chloride, creatinine, hematocrit, sodium). The discrepancy likely stems from different approaches to handling frequency ties, as the gold query uses DENSE_RANK which can return a different number of results than expected when multiple tests have the same frequency.